Show all the steps in clear way please and thanks in advance!
4. Solve for V, and I, in the circuit below. Answer in polar. 20 ΚΩ -j4 ΚΩ 10 ΚΩ Ο + > 2/30° V -j2 ΚΩ Το.
Answer #1

-jaka H ī loka volt to yo =V6 jakin 2 | 30° let assume opamp is ideal. ADL Copen loop gain) = 0 Va = V6 = V KCL at node Vi- VSubstitute v, in ega o 0 V, (0.15+j0.25) – 0.15 Vo = 0.15 /120° (0.15 +30.25)[0:45.730.5]vo -0.15 Vo = 0.5L1200 0.51120° 15 +